LT2 Technical Support Group

→ Europe/Zurich
532 Blackett Lab (Imperial College London)

532 Blackett Lab

Imperial College London

David Colling
Description
The meeting will be reachable by phone conference at 0845 862 0333 pin: 963852
Present: * IC: D. Colling, M. Aggarwal * IC-LESC: K. Sephton * QMUL: G.Mazza * UCL-HEP/CCC: Ben * BRUNEL/RHUL: Duncan * LHCB: Raja 1. Site Status * IC-HEP & IC-LESC -- MA: IC-HEP -- High load of biomed jobs on IC-HEP. Therefore IC-HEP site dropping out of BDII. Set limit per user job submission. -- KS: IC-LeSC -- New Viking cluster installed with SL4. Next step is to test the cluster on the production system. -- Need to know the current status of (sgm/prd) pool accounts from experiment point of view. -- Meeting with MA to discuss in detail. * UCL-HEP & UCL-CCC -- Some of UCL-CENTRAL cluster switched off over weekend for urgent air-conditioning repairs -- Number of nodes in UCL-CENTRAL has been reduced to make room for the new SRIF3-funded facility -- More jobs from LCG are required. William Hay reported a few jobs from (Atlas, biomed, lhcb and cms). CMS should install their software. * QMUL -- GM: SAM tests are improved. Still having problems with (SE free space) SAM test. -- GM: CMS jobs failing using CERN (RB). The problem is related to a specific node (fe04) used for CMS installation. Alex changed the firewall configuration of the machine. * RHUL/BRUNEL -- DR: DPM (Out of memory) problem with RHUL. The problem was previously observed by ATLAS users, it is new for CMS. Solution: Upgrade to most recent version of gLite. -- DR: Brunel upgraded their network to 300Mbits/s. But Iperf tests( from Glasgow) gives very low bandwidth ( 2Mb/s - 130 Mb/s). 2. Experiment problems/issues (If there any) * ATLAS -- none * CMS -- DC: CMS starting CSA07 from next week. FTS transfers (8 files) to London (Imperial, Brunel, RHUL). -- UCL, QMUL -- mainly MC production jobs. -- SW: Install CMS software at UCL-CC * LHCb -- Raja: No distributed analysis on Tier2's. Reconstruction at Tier 1's. Running test jobs on SL4 cpus. * Others: -- None 3. Open tickets & Action List Ongoing: 1730 Jobs from rb122.cern.ch failing at UKI-LT2-QMUL Ongoing: lcg utils problems (Imperial & RHUL) Closed: 1804 CE failure on ce02.esc.qmul.ac.uk (UKI-LT2-QMUL) -- All Actions: Open DC: -- Invite Lucio (QMUL) to join LT2 meetings. -- Dteam Meeting: The current status of (sgm/prd) pool accounts from experiment point of view. DR: -- set up LFC for LT2 VO, with help from Mona Lucio (QMUL): -- Follow up with Atlas production status and plans. SW: -- Install CMS software at UCL-CC All: -- Email site plans for SL4 upgrade -- Email quaterly report info to David 4. Who is going to join the dteam meeting tomorrow * David Colling & Duncan
There are minutes attached to this event. Show them.
    • 14:30 → 14:35
      Site Status 5m
      * IC-HEP & IC-LeSC * UCL-HEP && UCL-CC * QMUL * RHUL * BRUNEL
    • 14:35 → 14:40
      Open tickets & Action List 5m
    • 14:40 → 14:45
      Experiment problems/issues (If there any) 5m
      Atlas CMS LHCB Others
    • 14:45 → 14:47
      Dteam meeting representive this week 2m